Circular No. 7007

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1998du, 1998dv, 1998dw D. Reiss, Department of Astronomy, University of Washington, reports

that the Mount Stromlo Abell Cluster Supernova Search Team (cf. IAUC 6639) has detected three new apparent supernovae on V and R CCD images taken by A. Graham at the Mount Stromlo 1.27-m telescope (+ Macho Camera), with confirming images by S. Sabine at the same instrument on Aug. 28 (SNe 1998du and 1998dv) and Aug. 31 (SN 1998dw).

SN  1998 UT  α_{2000}  δ_{2000}  V  Offset
1998du Aug. 14 0^{h}45^{m}11^{s}.73 −63^{o}48′21′′.6 20.5 1′′.9 E, 0′′.8 N
1998dv Aug. 22 4 29 46.79 −61 30 25.4 20.2 2′′.1 W, 0′′.3 N
1998dw Aug. 28 1 09 11.56 −15 29 46.7 18.8 2′′.7 E, 1′′.1 N

SN 1998du is near Abell 2819 and was not present in images obtained with the same instrument on July 2, July 14, or Aug. 1 (limiting magnitudes V ∼ 21.0, 20.0, and 20.0, respectively). SN 1998dv is near Abell 3266 and was not visible in images obtained on Apr. 13, July 14, or Aug. 1 (limiting mag V ∼ 20.5, 20.0, and 20.5, respectively). The SN 1998dw is near Abell 151 and was not seen in images taken on 1997 Dec. 1 or 22 (limiting mag V ∼ 20.0 and 19.0, respectively).

XTE J0053–724 J. C. Lochner, Goddard Space Flight Center and Universities Space

Research Association, writes: “The transient x-ray system XTE J0053–724 (= 1WGA J0053.8–7226; cf. IAUC 6788, 6789, 6803) was detected in an observation by RXTE on Aug. 31. The measured 2–10-keV flux was estimated to be 3.0 × 10^{−11} erg s^{−1} cm^{−2}, with a photon index of 0.82 and a cutoff energy of 12.5 keV. Pulsations of 46.6 ± 0.1 s were observed, with a pulse fraction of ∼ 25 percent. This source was last detected in an observation on Apr. 8 while monitoring the second outburst of AX J0051–722 (IAUC 6858). This, along with the previous detection from late Nov. through mid-Dec. 1997, suggests a possible orbital period of this Be/neutron star system of ∼ 139 days. Continued optical monitoring is encouraged.”

V592 HERCULIS G. Masi, Ceccano, Italy, reports the following position for V592 Her

(cf. IAUC 7002): α = 16^{h}30^{m}56^{s}.42, δ = +21^{o}16′58′′.3 (equinox 2000.0; reference stars from USNO A1.0 catalogue).

1998 September 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green